5. The frequency distribution shows the salaries in million for specific year of top 25ceo’s in United States. Find the mean, variance and standard deviation. 

Class Boundaries Frequency 

15.5-20.5 13 

20.5-25.5   6 

25.5-30.5   4 

30.5-35.5 1 

35.5-40.5 1 


1
Expert's answer
2022-06-13T18:07:43-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 Classes & Frequency & Midpoint & M_i\cdot f_i \\
 & f_i & M_i \\
\hline
 15.5-20.5 & 13 & 18 & 234 \\
 \hdashline
 20.5-25.5 & 6 & 23 & 138 \\
 \hdashline
 25.5-30.5 & 4 & 28 & 112 \\
 \hdashline
 30.5-35.5 & 1 & 33 & 33 \\
 \hdashline
 35.5-40.5 & 1 & 38 & 38 \\
 \hdashline
\end{array}"

"\bar{x}=\dfrac{1}{N}\sum_iM_i\cdot f_i"

"=\dfrac{234+138+112+33+38}{13+6+4+1+1}=22.2"




"s^2=\dfrac{1}{N-1}\sum _i(M_i-\bar{x})^2\cdot f_i"

"=\dfrac{1}{25-1}((18-22.2)^2(13)+(23-22.2)^2(6)"

"+(28-22.2)^2(4)+(33-22.2)^2(1)"

"+(38-22.2)^2(1))=\dfrac{734}{24}\approx30.583333"

"s=\sqrt{s^2}=\sqrt{\dfrac{734}{24}}\approx5.530220"